answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

Is 'or' statement in probability mutually exclusive?

No. P(A or B) means A, B or (A and B).

For example, P(red card or a king) includes red kings.

What is probability of selecting a rational number from set of real number?

Zero, since there are infinitely more irrational numbers than rational numbers. Note that "zero probability" is not the same as "impossible" in this case. For more details, see the Wikipedia article on "Almost surely".

What type of probability used sample spaces to determine the numerical probability that an event will occur?

Discrete probability. It helps if the all the outcomes in the sample space are equally probable but that is not a necessity.

What is the largest value probability that can exist?

A probability must needs be a number between 0 and 1 (often expressed as 0% and 100%), inclusive.

Is there an actual strategy to winning Rock, Paper, Scissors?

Although technically Rock, Paper, Scissors is considered a game of chance, there are a few patterns to keep in mind that could give you a leg up on your competition.

Men are statistically more likely to throw rock first, whereas women are more likely to start with scissors. Further, if someone wins on a particular throw, they are more likely to start with the same choice in the next round. Similarly, if they lose, they’re likely to start throwing with the next item of the sequence. Ergo, if someone loses with rock, they’re most likely to start the next round with paper. Also, if a competitor throws the same sign consecutively, they are very unlikely to throw the same sign a third time, allowing you to rule out one possibility on the final throw.

Finally, making it a best 3 out of 5 is a good place to start improving your odds.

When a die is thrown what is the probability of getting not a composite number?

The composite numbers are 4 and 6; that is 2/6 of the numbers available. Therefore the odds on rolling such a number are "two out of 6" = 1/3 = 33.333%

The 3 in that answer is recurring.

Since the total of all probabilities which might occur in one event is 1.0 or 1 or 100%, it follows that the probability of not throwing a composite number is 4/6 = 2/3 = 66.66666666666666(recurring)%

What are some probability questions for a nine sided dice?

Q1: You can have a die and many dice, but what is "a dice"?

Q2: A 9-sided die cannot be a regular polyhedron. With an irregular polyhedron some outcomes are more likely than others. What is the shape?

Q3: Since there are 2606 different (non-isomorphic) 9-sided dice possible and each one will have a different probability distribution, how will you determine the probability distribution function for the faces on your die?

Q4: What are the numbers on the faces? Are they 1-9 or some other set, such as for a doubling die in backgammon?

Once you have got your answers to the above you can start thinking about the probability distribution function for the different numbers on the die. Then you can look at questions regarding multiple rolls.

How do you become a ballerina?

You have to lake lessons, studio or private lessons (recommended for best learning) You have to train HARD, about and hour a day if you want to be sucessful. Usually the ballerinas today started when they were very young (For best results start between the ages 3-8). You have to try your best to never take a break or miss a practice, everything counts! When you are older (18+) hopefully you have practiced Ballet for a minimum of 8 yrs. you must get discovered by signing up with a Ballet Company (example: American Ballet Theatre), you may be turned down but never lose hope, come back in a few months and try again, they will be convinced sooner or later! You will start with being in the "Corps de Ballet" Background dancers (Do some research for more info). The Ballerina is the main dancer and center of attention, so you have to go through about 2-3 levels in skill. Hope that answers your question, if you have a dream to become a Ballerina one day-- I hope you try your best!

What's one of the random number generators of a pair of dice?

If you just want to roll a die and not learn how to do it, you can type "roll a die" into Google, and it will have a six-sided die for you to roll right on that page. Other sizes of dice rollers will be available in the search results as well. If you want to learn how to do it in Excel though, that is below:

If you want to simulate the outcome of throwing a pair of dice (two die), it is easy and I will explain how to do this with Excel.

A single die has a discrete uniform distribution, with values from 1 to 6. Some programs have a built in function for computing random values. Excel does not, but it does have a built in continuous random number generator rand() from 0 to 1. Rand() will generate numbers above 0 and below 1, so if I multiply the numbers by 6, I will have numbers above 0 and below 6. I will take this number and use the int function to round down the number to the next lowest whole number, so now I have discrete numbers from 0 to 5. I can add one to these numbers and now I have random numbers 1 to 6.

So, in Excel you can use int(rand()*6)+1 to obtain the random numbers. I can put this equation in a second cell, and add them together. You can copy this equation many times and obtain many simulated throws of dice. In Excel, you can hit the F9 key and execute this program, obtaining different random numbers.

How can you put probability in a sentence?

I prefer to play the lotto games with the smaller prizes because the probability of winning something is much greater.

What is the variance of a standard deviation of 12.4?

Variance is std dev squared. Therefore, if std dev = 12.4, variance = 12.4^2 = 153.76.

Used to determine the probability of having specific offspring from two known parents?

Genotype is used to determine the probability of having specific offspring from two known parents.

What is the theoretical probability of tossing a coin?

That's a very complicated calculation, since it depends on so many variables that are

difficult to evaluate or quantify. A few of the factors that would influence it are:

-- whether or not you have a decision to make;

-- how you go about considering the pros and cons, and exactly when you decide to

leave the decision to chance, and to abide by the chance outcome;

-- whether or not you have access to a coin at that moment;

You would need to define your question a little more precisely before it would be possible to estimate an answer. For example, "what is the theoretical probability that any person anywhere on earth tosses a coin within the next one million years?" would give a very different estimated answer than if you asked, "what is the probability that a certain defined person tosses a coin within a defined five minute time span?"

Or your question could be interpreted as to mean, "What are the possible outcomes when tossing a coin, and what is the theoretical probability of each outcome?"

What is the defionition of probability?

Probability is the frequency that something in particular will happen, expressed in ratio with the frequency of anything happening. For instance, if you say that the probability of something is 0.25, then you can expect, over the long run, that repeated trials will result in 25% of the desired results and 75% of the non desired results.

A concrete example: In a standard deck of 52 cards, there are 12 face cards, the Jack, Queen, and King of Spades, Diamonds, Clubs, and Hearts. The probability of drawing a face card from a randomly shuffled deck is 12 in 52, or 3 in 13, or about 0.2308.

Probability is always a number between 0 and 1. A probability of 0 means that the desired event will never happen, and a probability of 1 means that the desired event will always happen. A consequence of this definition is that, if you say that the probability of something happening is p, then the probability of that something not happening is 1 - p.

How do you Quantum entangle?

The "easiest" way to do so is to cool a substance until it becomes a super-conductor. When that happens, you have Cooper Pairs, which are electrons with quantum entanglement.